OpenStudy (anonymous):

What is the slope of the line that passes through the points (4,2) and (10,5)?

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1) here x1=4;x2=10 y1=2; y2=5 now put the values ande check

OpenStudy (zale101):

\[m= \frac{ y_2-y_1}{ x_2-X_1 }\]

OpenStudy (anonymous):

but whats the answer?

OpenStudy (zale101):

Sorry, we don't give out answers. You lead you on finding the answer.

OpenStudy (zale101):

so do u get any of the formula i put above?

OpenStudy (zale101):

or what @madrockz said?

OpenStudy (anonymous):

nvm its 1/2

OpenStudy (anonymous):

(5-2)/(10-4) =3/6 =1/2

OpenStudy (zale101):

we lead you on finding the answer.*

OpenStudy (zale101):

@madrockz is correct :D

OpenStudy (dan815):